OpenStudy (anonymous):

The width of a rectangle is two-thirds of its length, and its area is 216 square meters. What is the length? 12 m 18 m 27 m

Parth (parthkohli):

\[\Large x * {2x \over 3} = 216 \] Solve it.

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

L = L w = (2/3)L A = 216 L x w = A Substitute A and w, then solve for L L x (2/3)L = 216 (2/3)L^2=216 L^2=324 L=18, therefore w=(2/3)L=(2/3)18=12

Parth (parthkohli):

\(\Large x^{2} = 324\) \(\Large x = 18\) Width is 12, as 2x/3 = 12